OpenStudy (anonymous):

√(x+4) -2/x=1/√(x+4) +2 How do you simplify this?

OpenStudy (joannablackwelder):

Does it look like this?:\[\sqrt{x+4}-\frac{ 2 }{ x }=\frac{ 1 }{ \sqrt{x=4} }+2\]

OpenStudy (joannablackwelder):

Should be x+4 in the denominator on the right

OpenStudy (anonymous):